1 Help with Check List Items

1.1 Help with Page 1

What Where to find it

How to do it

DATE (month-day-year)  As given by Run Control  Enter as MM/DD/YY
Time (24 hour clock)  As given by Run Control  Write the time into the cell

Your Name

As given by your parents Write legibly

Beam Conditions from Overview Screen, DAQ, MEDMs, and Scalers Beam Energy (MeV) Large TV monitor on top of console,

displaying accelerator program.

Get the beam energy of Hall C and write it down

Beam Current Look for the BCM screen on slow control monitor. The two values indicate two different measurements and should be similar.

Write the beam current in the cell.

IPM3H00AX/AY nominal? (medm)

Xterm on the fartop- left. Look for

the BPM screen.

Record the displayed beam position. Compare it with the nominal position in latest run plan.

IPM3H00BX/BY nominal? (medm)

Xterm on the fartop- left. Look for

the BPM screen.

Record the displayed beam position. Compare it with the nominal position in latest run plan.

SLow raster (circular?) Check oscilliscope labelled “Slow raster Display” in electronics room

If you see a circular pattern write YES

Fast Raster R_{max},Ix,Iy (medm)

Xterm on the top of the right console. Look for the raster hlc FR window.

In the “set spiral magnet current” field, record: R_{max}= Radius(mm) Ix=MRAT301H Iy=MRAT301V

Scalers from the bigcal.tcl display on “Slow Cnntrols” monitor Run Number See the RunControl window for the current run. Record items below for a recently taken and analyzed run.

Enter the run number which gives the results noted in the next few lines.

HMS Trigger Rate Read the rates from scaler running on

slow control monitor.

BETA1 Trigger Rate Read the rates from scaler running on

slow control monitor.

BETA2 Trigger Rate Read the rates from scaler running on slow control monitor.

Computer Deadtime Read the deadtime from scaler running on slow control monitor.

Record the Computer deadtime.

Walk-through of Electronics Room WC Thresholds HMS Rack CH03B10. Three digital voltage

readouts. Should be 5.5 v for HMS

chambers.

Unser Monitor Temperature top of rack CH03B12, LED digits.

Nominal 110 degrees.

Record the reading.

Gas-Yard Checks (Gas Yard Key is #29 Argon Bottle High/Low (psig) Outside gas shed, under roof, right-

center.

Record the two pressure gauge readings on the argon gas manifold.

Ethane Bottle High/Low (psig)

Outside gas shed, under roof, on the right.

Record the two pressures on the ethane gas manifold.

Helium bottle Pressure Outside gas shed, under roof, on the left.

Record the bottle pressure and the flowmeter reading.

Helium bottle Flow Outside gas shed, under roof, on the left.

Record the bottle pressure and the flowmeter reading.

Mixer Flows: Argon=CH1 Ethane=CH2 Inside gas shed, on the CRT display - channels 1 and 2. (Press ONLY “esc” to wake display.)

Record the values on the “Actual flow” line. The two values should be about equal for a 50/50 mixture.

Mixer Outlet Pressure Inside gas shed, Pressure Gauge on left-hand rack. Read “psi” scale.

Record the pressure. Should be ≈ 9.5psi

Alcohol Temperature (^{∘} C) Inside gas shed, digital temperature display in right-hand rack.

Record the temperature.

Controller Temperature (^{∘} C) Inside gas shed, digital temperature display in right-hand rack.

Record the temperature.

Alcohol Bubbling? (HMS and FPP)

Inside gas shed, within refrigerator (right=HMS).

QUICKLY open/close door, verify that gas is bubbling through the alcohol inside the glass dome.

1.2 Help with Page 2

What Where to find it

How to do it

DATE (month-day-year)  See what you wrote on page 1.  Enter it here, too.
Time (24 hour clock)  See what you wrote on page 1.  Enter it here, too.

HMS slit telnet hctsv7 2003 (HMS) or hctsv5 2005 (SOS). See instructions posted above “SLOW CONTROLS” Xterm or see the Collimators Howto [?].

Determine and record the slit position as either PION or SIEVE or SMALL or LARGE.

Scattering Chamber Vacuum Target cryo screen. Ask target operator if needed.

Record the reading.

Target

Ask target operator. Record which target is in the beam.

HMS Shutter Open? See the LEDs labeled “HMS SHUT-TER” located above the “DATA AC-QUISITION” terminal.

Shutter must be “OUT” when taking beam.

HMS angle (TV/GUI) TV Monitor 4 and Spectrometer Rotation GUI Read the angle and write it down. TV and GUI should agree.

Hall Survey (TV) Scan all the cameras in the hall looking for anything unusual.

Verify that you performed the inspection by checking off the box.

HMS Cer. Temp/Press (TV) Two LED readouts visible on TV monitor #7. Select the camera labeled “HMS CERENK”.

Record the values.

High Voltages: HV ON and GUI Alarm Enabled? All HMS Detectors HMS HV GUI on cdaqs3 xterm. Verify that all GUI is “RUNNING”

and not indicating any bad channels.

All BETA Detectors BETA HV GUI on cdaqs3 xterm. Verify that all GUI is “RUNNING”

and not indicating any bad channels.

All BigCal Detectors BigCal HV is on cdaqs3 xterm and has three GUIS ‘GEP HV CONTROLS, TEMP HV CONTROLS, and High Voltage System Control.

Verify that all three GUIs are “RUN-NING” and not indicating any bad channels.

Spectrometer Magnets HMS Q1,2,3 and D Currents: ± setpt PC Monitor in center of control console. See HMS magnet controls Howto [?] for how to navigate.

Write down current polarities and setpoints.

HMS Q1, Q2, Q3 Hall Fields: ± reading On TV monitor #5. Button labeled HMS NMR (it’s really a Hall probe). Top to bottom, the three numbers are Hall Probe readings for HMS Q1, Q2, Q3.

Write down the three numbers including sign.

HMS D NMR Field: ± readback PC Monitor in center of control console. See HMS magnet controls Howto [?] for how to navigate.

Write down the NMR readout.

HMS Q1Q2, Q3, D N2/He Level PC Monitor in center of control console. See HMS magnet controls Howto [?] for how to navigate.

Record He and N2 levels from their respective screens.

Enter Screen Snapshots in hclog of BPM medm screen , BCM medm screen Start an hclog entry. Use the

                                  Grab Screen and Grab 2nd Picture

to add the required displays to the log entry.

Scan this sheet into hclog when full.

Any Xterm booted on cdaqh1 Place the sheet to be scanned on the scanner in rack CH03B08. Start an hclog entry. Use the ScanIt! button to scan the page, then use the Grab Screen button to grab the display of the scan.